# ARTIST: Improving the Generation of Text-rich Images with Disentangled Diffusion Models

### v2

@@ Abstract
Diffusion models have demonstrated exceptional capabilities in generating a broad spectrum of visual content, yet their proficiency in rendering text is still limited: they often generate inaccurate characters or words that fail to blend well with the underlying image. To address these shortcomings, we introduce a new framework named ARTIST. This framework incorporates a dedicated textual diffusion model to specifically focus on the learning of text structures. Initially, we pretrain this textual model to capture the intricacies of text representation. Subsequently, we finetune a visual diffusion model, enabling it to assimilate textual structure information from the pretrained textual model. This disentangled architecture design and the training strategy significantly enhance the text rendering ability of the diffusion models for text-rich image generation. Additionally, we leverage the capabilities of pretrained large language models to better interpret user intentions, contributing to improved generation quality. Empirical results on the MARIO-Eval benchmark underscore the effectiveness of the proposed method, showing an improvement of up to 15% in various metrics.
@@ 1 Introduction
The field of text-to-image generation has made remarkable progress, especially with the rise of diffusion models [10, 28, 26, 31, 45, 51, 30]. These models not only have demonstrated prowess in generating high-fidelity images, but have also showcased versatile applications, such as image inpainting [19, 7, 13] , denoising [43, 44], video generation [8, 2], and style transfer [40, 50]. Along with these advancements, the ability to generate images containing legible text, another realm of significance, is becoming increasingly worthy of attention. Specifically, this need is evident in everyday scenarios, where images with text are commonplace, from advertisements to road signs, posters, and book covers. Crafting these text-rich images manually demands skilled expertise and a significant time commitment. However, a significant shortcoming lies in the current state-of-the-art diffusion-based generative models; they often render text portions that are virtually unreadable, akin to gibberish, undermining the aesthetic and functional value of the generated images.
@@
Figure 1: Generated examples from our ARTIST. It adeptly identifies the text intended to be generated in the image from the given prompts, regardless of explicitly marked by quotes. The generated text is legible and complements the visual elements, enhancing the overall coherence of the design.
@@
To tackle the challenges posed by the quality of text generated by diffusion models, two primary pathways have been explored. Firstly, the traditional approach involves leveraging image-editing tools to superimpose text onto images directly. However, this frequently introduces unnatural artifacts, especially when dealing with intricate textures or varying lighting conditions in the background of the image. In contrast, recent research efforts aim to refine the diffusion models themselves for improved text quality. For example, recent innovations such as Imagen [31], eDiff-I [1], and DeepFloyd [35] discovered that the use of T5 series text encoders [25] led to better results compared to using the CLIP text encoder [23]. Similarly, Liu et al.[17] integrated character-aware text encoders to enhance the quality of text rendering. However, these advancements primarily revolve around optimizing text encoders and do not necessarily grant more control over the holistic generation process. In parallel, GlyphDraw [20] has enhanced model controllability by conditionally focusing on the positioning and architecture of Chinese characters. Still, its utility remains limited, as it cannot cater to scenarios that require multiple text bounding boxes, making it less suitable for prevalent text-image formats like posters and book covers. TextDiffuser [6] represents a recent advancement in the realm of enhancing the quality of text in generated images. While it undeniably marks a significant step forward, it is not without limitations. One of the prominent challenges is the model’s dependence on manual efforts when it comes to recognizing key terms from prompts, making the process less efficient than desired. Additionally, even though TextDiffuser exhibits improved capabilities, the accuracy rate of its Optical Character Recognition (OCR) evaluation on generated text still leaves room for optimization.
@@
In this paper, our primary objective is to develop a more efficient system for text-to-image generation. This system would eliminate the need for manual post-production adjustments, ensuring superior text quality within generated images. Upon meticulous investigation, we have identified the primary challenges we aim to address. Firstly, automation of accurately discerning which words from a provided text prompt should be incorporated into the image. In platforms such as TextDiffuser, this identification requires human involvement, typically by highlighting specific terms with quotation marks, diminishing the platform’s efficiency and automation. Second, there is the challenge of adeptly generating images that seamlessly integrate top-quality text, ensuring adherence to a predetermined layout, and emphasizing the text in the generated images. To address the aforementioned challenges, we turned to the latest advancements in large language models (LLMs) [22, 36, 37]. These models have showcased outstanding expertise in understanding and processing intricate linguistic patterns, making them perfectly suited for our needs. Inspired by their prowess, we formulated a strategy that employs LLMs to surmount the initial hurdle of pinpointing the keywords. To tackle the subsequent challenge, we introduce a novel two-stage approach named ARTIST, detailed in Figure 2. Specifically, ARTIST’s dual stages are dedicated to mastering text structure and refining visual aesthetics in that order. The moniker “ARTIST" encapsulates the essence of the Ability of Rendering Text can be Improved by diSentanglemenT. By merging our method with LLMs’ capabilities, we have achieved a marked improvement in the quality of text embedded within images.
@@ 2.1 Text-to-image Generation
2.2 Text Rendering & Optical Character Recognition
Despite the rapid development of diffusion-based Text-to-image Generation, existing methodologies still struggle with generating precise and consistent textual renderings. Various approaches, like Imagen [31], eDiff-I [1], and DeepFolyd [35], have leveraged the prowess of expansive language models (notably large T5 [25]) to bolster their textual accuracy. The study in [17] highlights a limitation in which traditional text encoders overlook token length, prompting them to propose a character-sensitive alternative. Simultaneously, GlyphDraw [20] focuses on producing superior images integrated with Chinese texts, guided by textual positioning and glyph imagery. GlyphControl [46] further enhances this approach by adjusting the text alignment according to its location, implicitly incorporating elements like font size and text box positioning. A recent study, AnyText [38], utilizes a diffusion pipeline with an auxiliary latent module and a text embedding module to improve the text generation, editing, and integration with the image background. The techniques of Textdiffuser [6] harnesses the Transformer model to discern keyword layouts, promoting multiline text generation. Then, it further employs character-based segmentation masks as a prior, offering flexibility in control to cater to user specifications. However, it still depends on manual efforts when it comes to recognizing key terms from prompts. Although Textdiffuser-2 [5] employs LLMs to enhance the interpretation of prompts, the improvements in the quality of generated images with text remain modest. This indicates that although the system has improved in understanding user inputs, converting these advancements into better visual results still demands additional optimization. Hence, we believe it is necessary to develop a more effective framework that captures the intricacies of text representation and then seamlessly integrates textual structures into images, aiming for more coherent visual outputs. From an evaluative standpoint, TextDiffuser introduced the benchmark called MARIO-Eval. This was accomplished by collecting and synthesizing insights from previous works, including the DrawBench from Imagen [31], the DrawTextCreative by [17], among others.
@@ 3 Methodology
  • It is challenging to simultaneously learn visual appearance and text structure with a single model;
  • Existing datasets cannot cover all the words and their possible combinations, making it hard to learn text structure from these limited noisy data.
In this work, we propose to mitigate the aforementioned challenges by utilizing separate modules to learn text structure and visual appearance. Furthermore, these two modules are trained separately, making it possible to learn text structure with synthetic data constructed by ourselves, which also tackles the problem of limited data. To ensure disentanglement and prevent information leakage, our text module and visual module take different prompts as inputs. However, it can be inefficient and user-hostile if these prompts have to be manually designed by the user themselves at inference. Fortunately, because of the recent success of large-language models (LLMs) [3, 22, 36, 37], we propose to utilize pretrained LLMs to infer user’s intention, provide accurate prompts for both modules. With the help of LLM, user’s input can be either precise or vague, leading to better interactive experience.
@@ 3.1 LLM-based Prompt Understanding
TextDiffuser [6] proposes to train a transformer model to extract texts which are expected to be shown on generated images. However, there appears to be a limitation associated with this method: their model is only capable of detecting words enclosed by quotation marks. This happens because most keywords from the training dataset are enclosed by quotation marks inside the captions. A model trained on these samples will lack generalization ability or even be overfitting. Moreover, their model even fails to generate the desired image according to the prompt “a poster of Batman with word Batman on it", as it does not understand that the word “Batman" should be presented on the image. On the contrary, we propose to provide pretrained LLM with an open-domain prompt and rely on its capability to identify the essentials autonomously. Because of its vast open-domain knowledge, LLM is able to understand user intents better and thus can generalize to complicated scenarios. Given vague prompts, LLM can discern and propose which words or text elements to incorporate, resulting in more coherent and aesthetically appealing suggestions.
@@ 3.2 Learning Text Structure
A diffusion model, denoted as a text module, is introduced to learn text structures. Specifically, this text module is trained to take bounding boxes as input and generates a black-white image with only text on it. Because we do not require it to learn any visual effects, the prompt for this module is the word to be rendered, and the dataset can be simply constructed using standard rendering libraries Specifically, we construct two large-scale datasets to train this module, as described below.
@@
  • Word-level dataset: Our first dataset consists of around 10 million black-white images with only a single word on it. Specifically, to construct each data sample, we first randomly select a word from the vocabulary of the CLIP text encoder, then render this word with a random font and size on a black image. Meanwhile, we can also obtain the ground-truth bounding box for each rendered word. This dataset will be further augmented during training by randomly moving both the word and bounding box to a new position, resulting in an infinite number of effective samples;
@@
  • Sentence-level dataset: Although our word-level dataset contains massive structure information of English words, it contains no layout information about how these words should be placed and combined on the target image. To this end, we construct our second dataset which contains 50 millions of black-white images by utilizing MARIO-10M dataset [6]. Specifically, we use the ground-truth text and layout information from MARIO-10M samples and render the same text with randomly selected fonts on black images following the same layout.
@@
After constructing both datasets, we train the text module in two stages, inside the latent space of a VAE [14] following Rombach et al. [28]. In the first stage, the diffusion model is trained to take a single bounding box and a target word as inputs then generate an black-white image with the word rendered on it. Then, this module will be finetuned on our sentence-level dataset as the second stage so that it can take multiple bounding boxes and multiple s as inputs. Our s are denoted as $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{n}\}$, composed of the words to be generated in the image. Our input mask is an image containing the corresponding bounding boxes $\{m_{1},m_{2},\ldots m_{n}\}$ that indicate the position of each word. During training, the original text-only image and input mask image are first encoded into latent space features $\mathbf{z}$ and $\mathbf{m}$. Then we sample a time step tsimilar-$t\sim$ Uniform $\left(0,T\right)$ and a Gaussian noise $\boldsymbol{\epsilon}$ to corrupt the original feature, yielding $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ where $\overline{\alpha}_{T}$ is the coefficient of the diffusion process introduced in [10]. $\mathbf{z}_{t}$ and $\mathbf{m}$ are concatenated in the feature channel as input for the diffusion model, which will be trained with the diffusion loss between the sampled $\boldsymbol{\epsilon}$ and the predicated noise $\boldsymbol{\epsilon}_{\theta}$:
@@
Some generated examples are shown in Figure 3, which illustrates that our text module can successfully generate a target image with desired texts on it in different styles.
In our implementation, MARIO-10M dataset [6] is used to train this module following previous work for a fair comparison. $\overline{\mathcal{P}},\mathcal{P},\mathbf{m}$ have already been prepared in MARIO-10M. At inference time, LLM will be utilized to infer $\overline{\mathcal{P}},\mathcal{P},\mathbf{m}$ as mentioned in Section 3.1.
@@ 3.3 Learning Visual Appearance
During the training of visual module, the text module will be frozen, and only the newly-introduced layers and visual module will be finetuned. Differently from the text module described in Section 3.2, whose training text only contains the target words to be rendered, the training data for our visual module has to contain visual descriptions of the image so that the model can successfully learn to generate visual contents based on the user’s input.
Let $\overline{\mathcal{P}}$ be the prompt which contains visual descriptions of the image, $\mathcal{P}$ be the s as defined in Section 3.2. We feed $\mathcal{P}$ and the input mask $\mathbf{m}$ as mentioned in Section 3.2 into the pretrained text module. The intermediate features from the text module denoted as $\{f_{i}(\mathbf{z}_{t},\mathbf{m},\mathcal{P},t)\}_{i=1}^{k}$ will be injected into the visual module, where $k$ stands for the number of intermediate features. The visual module will be trained with diffusion loss:
$\mathcal{L}_{\text{visual}}=\mathbb{E}\left[\left\|\epsilon-\epsilon_{\theta}\left(x_{t},\{f_{i}(z_{t},m,\mathcal{P},t)\}_{i=1}^{k},m,\overline{\mathcal{P}},t\right)\right\|_{2}^{2}\right]\vspace{-3mm}$

### v3

@@ Abstract
Diffusion models have demonstrated exceptional capabilities in generating a broad spectrum of visual content, yet their proficiency in rendering text is still limited: they often generate inaccurate characters or words that fail to blend well with the underlying image. To address these shortcomings, we introduce a novel framework named ARTIST, which incorporates a dedicated textual diffusion model to focus on the learning of text structures specifically. Initially, we pretrain this textual model to capture the intricacies of text representation. Subsequently, we finetune a visual diffusion model, enabling it to assimilate textual structure information from the pretrained textual model. This disentangled architecture design and training strategy significantly enhance the text rendering ability of the diffusion models for text-rich image generation. Additionally, we leverage the capabilities of pretrained large language models to interpret user intentions better, contributing to improved generation quality. Empirical results on the MARIO-Eval benchmark underscore the effectiveness of the proposed method, showing an improvement of up to 15% in various metrics.
@@ 1 Introduction
The field of text-to-image generation has made remarkable progress, especially with the rise of diffusion models [13, 35, 32, 37, 53, 74, 36]. These models not only have demonstrated prowess in generating high-fidelity images, but have also showcased versatile applications, such as image inpainting [25, 9, 17] , denoising [51, 52], video generation [10, 2], and style transfer [48, 72]. Along with these advancements, the ability to generate images containing legible text, another realm of significance, is becoming increasingly worthy of attention. Specifically, this need is evident in everyday scenarios, where images with text are commonplace, from advertisements to road signs, posters, and book covers. Crafting these text-rich images manually demands skilled expertise and a significant time commitment. However, a significant shortcoming lies in the current state-of-the-art diffusion-based generative models; they often render text portions that are virtually unreadable, akin to gibberish, undermining the aesthetic and functional value of the generated images. Therefore, if generative AI, powered by diffusion models, can produce such images, it could revolutionize design workflows, inspire creativity, and alleviate designers’ workload.
@@
Figure 1: Generated examples from our ARTIST. Our framework adeptly identifies the text intended to be generated in the image from the given prompts, regardless of explicit marking by quotes. The generated text is legible and complements the visual elements, enhancing the overall coherence of the design.
@@
To tackle the challenges posed by the quality of text generated by diffusion models, two primary pathways have been explored. Firstly, the traditional approach involves leveraging image-editing tools to superimpose text onto images directly. However, this frequently introduces unnatural artifacts, especially when dealing with intricate textures or varying lighting conditions in the background of the image. In contrast, recent research efforts aim to refine the diffusion models themselves for improved text quality. For example, recent innovations such as Imagen [37], eDiff-I [1], and DeepFloyd [41] discovered that the use of T5 series text encoders [31] led to better results compared to using the CLIP text encoder [29]. Similarly, Liu et al.[23] integrated character-aware text encoders to enhance the quality of text rendering. However, these advancements primarily revolve around optimizing text encoders and do not necessarily grant more control over the holistic generation process. On a parallel track, GlyphDraw [26] has enhanced model controllability by conditionally focusing on the positioning and architecture of Chinese characters. Still, its utility remains limited, as it cannot cater to scenarios that require multiple text bounding boxes, making it less suitable for prevalent text-image formats like posters and book covers. TextDiffuser [7] represents a recent advancement in the realm of enhancing the quality of text in generated images. While it undeniably marks a significant step forward, it is not without limitations. One of the prominent challenges is the model’s dependence on manual efforts when it comes to recognizing key terms from prompts, making the process less efficient than desired. Additionally, even though TextDiffuser exhibits improved capabilities, the accuracy rate of its Optical Character Recognition (OCR) evaluation on generated text still leaves room for optimization.
@@
In this paper, our primary objective is to develop a more efficient system for text-to-image generation. This system would eliminate the need for manual post-production adjustments, ensuring superior text quality within generated images. Upon meticulous investigation, we have identified the primary challenges we aim to address. Firstly, automation of accurately discerning which words from a provided text prompt should be incorporated into the image. In platforms such as TextDiffuser, this identification requires human involvement, typically by highlighting specific terms with quotation marks, diminishing the platform’s efficiency and automation. Second, there is the challenge of adeptly generating images that seamlessly integrate top-quality text, ensuring adherence to a predetermined layout, and emphasizing in the generated images.
To address the aforementioned challenges, we turned to the latest advancements in natural language processing and large language models (LLMs) [28, 43, 47, 18, 57, 59, 20, 63]. LLMs have showcased outstanding expertise in understanding and processing intricate linguistic patterns, making them perfectly suited for our needs. Inspired by their prowess, we formulated a strategy that employs large language models to surmount the initial hurdle of pinpointing the keywords. To tackle the subsequent challenge, we introduce a novel two-stage approach named ARTIST, detailed in Figure 2. Specifically, ARTIST’s dual stages are dedicated to mastering text structure and refining visual aesthetics in that order. The moniker “ARTIST" encapsulates the essence of the Ability of Rendering Text can be Improved by diSentanglemenT. By merging our proposed methodology with LLMs’ capabilities, we have achieved a marked improvement in the quality of text embedded within images.
@@ 2.1 Text-to-image Generation
2.2 Text Rendering & OCR
Despite the rapid development of diffusion-based Text-to-image Generation, existing methodologies still struggle with generating precise and consistent textual renderings. Various approaches, like Imagen [37], eDiff-I [1], and DeepFolyd [41], have leveraged the prowess of expansive language models (notably large T5 [31]) to bolster their textual accuracy. The study in [23] highlights a limitation in which traditional text encoders overlook token length, prompting them to propose a character-sensitive alternative. Simultaneously, GlyphDraw [26] focuses on producing superior images integrated with Chinese texts, guided by textual positioning and glyph imagery. GlyphControl [56] further enhances this approach by adjusting the text alignment according to its location, implicitly incorporating elements like font size and text box positioning. A recent study, AnyText [45], utilizes a diffusion pipeline with an auxiliary latent module and a text embedding module to improve the text generation, editing, and integration with the image background. The techniques of Textdiffuser [7] harnesses the Transformer model to discern keyword layouts, promoting multiline text generation. Then, it further employs character-based segmentation masks as a prior, offering flexibility in control to cater to user specifications. However, it still depends on manual efforts when it comes to recognizing key terms from prompts. Although Textdiffuser-2 [6] employs LLMs to enhance the interpretation of prompts, the improvements in the quality of generated images with text remain modest. This indicates that although the system has improved in understanding user inputs, converting these advancements into better visual results still demands additional optimization. Hence, we believe it is necessary to develop a more effective framework that captures the intricacies of text representation and seamlessly integrates textual structures into images, aiming for more coherent outputs.
@@ 3 Methodology
  • It is challenging to simultaneously learn visual appearance and text structure with single model;
  • Existing datasets are not able to cover all the words and their possible combinations, making it hard to learn text structure from these limited noisy data.
In this work, we propose to mitigate the aforementioned challenges by utilizing separate modules to learn text structure and visual appearance. Furthermore, these two modules are trained separately, making it possible to learn text structure with synthetic data constructed by ourselves, which also tackles the problem of limited data. To ensure disentanglement and prevent information leakage, our text module and visual module take different prompts as inputs. However, it can be inefficient and user-hostile if these prompts have to be manually designed by the user themselves at inference. Fortunately, because of the recent success of large-language models (LLMs) [3, 44, 65, 61, 64, 66], we propose to utilize pre-trained LLMs to infer user’s intention, provide accurate prompts for both modules. With the help of LLM, user’s input can be either precise or vague, leading to a better interactive experience. Due to the disentangled nature of our framework, it is also friendly for future implementations that leverage distributed training [62, 54, 8, 42, 16, 11, 60, 55] to utilize computational resources while safeguarding data privacy.
@@ 3.1 LLM-based Prompt Understanding
TextDiffuser [7] proposes to train a transformer model to extract texts that are expected to be shown on generated images. However, there appears to be a limitation associated with this method: their model is only capable of detecting words enclosed by quotation marks. This happens because most keywords from the training dataset are enclosed by quotation marks inside the captions. A model trained on these samples will lack generalization ability or even be overfitting. Moreover, their model even fails to generate the desired image according to the prompt “a poster of Batman with the word Batman on it", as it does not understand that the word “Batman" should be presented on the image.
On the contrary, we propose to provide pre-trained LLM with an open-domain prompt and rely on its capability to autonomously identify the essentials. Because of its vast open-domain knowledge, LLM is able to better understand user intents and thus can generalize to complicated scenarios. Given vague prompts, LLM is able to discern and propose which words or text elements to incorporate, resulting in more coherent and aesthetically appealing suggestions.
@@ 3.2 Learning Text Structure
A diffusion model, denoted as a text module, is introduced to learn text structures. Specifically, this text module is trained to take bounding boxes as input and generates a black-white image with only text on it. From Bayesian learning [68, 67, 71, 73, 5] perspective, our text module essentially learns a more reliable prior distribution for the subsequent generation process. Because we do not require it to learn any visual effects, the prompt for this module is the word to be rendered and the dataset can be simply constructed using standard rendering libraries Specifically, we construct two large-scale datasets to train this module, as described below.
Word-level dataset:
@@
Our first dataset consists of around 10 millions of black-white images with only single word on them. Specifically, to construct each data sample, we first randomly select a word from the vocabulary of CLIP text encoder, then render this word with random font and size on a black image. Meanwhile, we can also obtain the ground-truth bounding box for each rendered word. This dataset will be further augmented during training by randomly moving both the word and bounding box to a new position, resulting in more effective samples;
Sentence-level dataset:
@@
Although our word-level dataset contains massive structure information of English words, it contains no layout information about how should these words be placed and combined on the target image. To this end, we construct our second dataset which contains 50 millions of black-white images by utilizing MARIO-10M dataset [7]. Specifically, we use the ground-truth text and layout information from MARIO-10M samples, and render the same text with randomly selected fonts on black images following the same layout.
@@
After constructing both datasets, we train the text module in two stages, inside the latent space of a VAE [19] following Rombach et al. [35]. In the first stage, the diffusion model is trained to take a single bounding box and a target word as inputs, then generate a black-white image with the word rendered on it. Then, this module will be fine-tuned on our sentence-level dataset as the second stage so that it can take multiple bounding boxes and multiple s as inputs. Our s are denoted as $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{n}\}$, composed of the words to be generated in the image. Our input mask is an image containing the corresponding bounding boxes $\{m_{1},m_{2},\ldots m_{n}\}$ that indicate the position of each word. During training, the original text-only image and input mask image are first encoded into latent space features $\mathbf{z}$ and $\mathbf{m}$. Then we sample a time step tsimilar-$t\sim$ Uniform $\left(0,T\right)$ and a Gaussian noise $\boldsymbol{\epsilon}$ to corrupt the original feature, yielding $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ where $\overline{\alpha}_{T}$ is the coefficient of the diffusion process introduced in [13]. $\mathbf{z}_{t}$ and $\mathbf{m}$ are concatenated in the feature channel as input for diffusion model, which will be trained with the diffusion loss between the sampled $\boldsymbol{\epsilon}$ and the predicated noise $\boldsymbol{\epsilon}_{\theta}$:
@@
Some generated examples are shown in Figure 3, which illustrates that our text module can successfully generate a target image with desired texts on it in different styles. In our implementation, MARIO-10M dataset [7] is used to train this module following previous work for a fair comparison. $\overline{\mathcal{P}},\mathcal{P},\mathbf{m}$ have already been prepared in MARIO-10M. At inference time, LLM will be utilized to infer $\overline{\mathcal{P}},\mathcal{P},\mathbf{m}$ as mentioned in Section 3.1.
@@ 3.3 Learning Visual Appearance
During the training of the visual module, the text module will be frozen, and only the newly introduced layers and visual module will be fine-tuned. Differently from the text module described in Section 3.2, whose training text only contains the target words to be rendered, the training data for our visual module has to contain visual descriptions of the image so that the model can successfully learn to generate visual contents based on the user’s input.
Let $\overline{\mathcal{P}}$ be the prompt which contains visual descriptions of the image, $\mathcal{P}$ be the s as defined in Section 3.2. We feed $\mathcal{P}$ and the input mask $\mathbf{m}$ as mentioned in Section 3.2 into the pre-trained text module. The intermediate features from the text module denoted as $\{f_{i}(\mathbf{z}_{t},\mathbf{m},\mathcal{P},t)\}_{i=1}^{k}$ will be injected into the visual module, where $k$ stands for the number of intermediate features. The visual module will be trained with diffusion loss:
$\vspace{-4mm}\mathcal{L}_{\text{visual}}=\mathbb{E}\left[\left\|\epsilon-\epsilon_{\theta}\left(x_{t},\{f_{i}(z_{t},m,\mathcal{P},t)\}_{i=1}^{k},m,\overline{\mathcal{P}},t\right)\right\|_{2}^{2}\right]\vspace{-6mm}$